Aksana and Ebony are competing in a wheelchair walkathon. The distance Aksana has traveled with respect to time is represented by the equation d=0.21t. In the equation, d is the distance in miles and t is the time in minutes. The distance Ebony has traveled with respect to the time is represented in the table below. If they continue at their current speeds, who will finish the marathon sooner?
Ebony's Race Data
Time(min) Distance(mi)
5 0.85
10 1.7
15 2.55

Answer by Theo(13342)

formula for aksana is d = .21t

if aksana travels for 5 minutes, her distance is .21 * 5 = 1.05 miles.
compare this to ebony's distance of .85 miles, and aksana is ahead.

if aksana travels for 10 minutes, her distance is .21 * 10 = 2.1 miles.
compare this to ebony's distance of 1.7 miles and aksana is still ahead.

if aksana travels for 15 minutes, her distance is .21 * 15 = 3.15 miles.
compare this to ebony's distance of 3.55 miles, and aksana is still ahead.

it appears that aksana will win the race.

you can calculate ebony's rate for 5, 10, and 15 minutes, to see if she is consistent.

the formula to use is d = rt.

solve for r to get r = d/t.

at 5 minutes and .85 miles, you get r = .85/5 = .17
at 10 minutes and 1.7 miles, you get r = 1.7/10 = .17
at 15 minutes and 2.55 miles, you get r = 2.55/15 = .17

ebony's rate is consistent at .17 miles per minute.

aksana's rate of travel is .21 miles per minute.
ebony's rate of travel is .17 miles per minute.

at these rates, aksana will win the marathon.